# Challenging Assumptions in Learning Generic Text Style Embeddings

### v1

@@ 1 Introduction
Learning meaningful representations of text has received much attention recently as models pre-trained on large corpora have become the standard for extracting representations that capture prior knowledge (i.a. Devlin et al., 2019; Radford and Narasimhan, 2018). However, most pre-trained models focus on general aspects, as they are trained with standard or masked language modeling objectives. Therefore, they are suboptimal for NLP tasks focusing on a specific aspect, such as style. There are many notable NLP tasks, such as identifying chatbot-written text based on its style (e.g., Soto et al. (2024)) or style transfer where models rely on dedicated style representations (e.g., John et al. (2019)).
@@
Based on the hypothesis that low-level style changes compose high-level style changes, this work explores learning generic, sentence-level style representations. We take a pre-trained encoding function producing general-purpose representations and fine-tune it to distinguish between low-level stylistic changes using contrastive learning and cross-entropy loss. We hypothesize that the resulting text encoder generalizes to high-level styles, applying the view by Lyu et al. (2021) on how low-level stylistic changes compose high-level changes to representation learning for styles.
@@ 3 Method
We combine this assumption with previous insights on how contrastive learning supports learning generic style embeddings (Patel et al., 2024) to learn generic style representations from low-level stylistic changes that generalize to composed, high-level styles.
@@ 4.1 Experimental Setup
Data
We conducted our experiments on datasets that take a data-driven approach, containing high-level stylistic changes, to define a text style. These are used in many recent works on text style transfer. We experiment with the Bible corpus Carlson et al. (2018) with eight different styles of the Bible (249K sentences), Grammarly’s Yahoo Answers Formality Corpus (GYAFC) Rao and Tetreault (2018) (113K sentences) in two styles, a collection of Shakespeare plays in Shakespearean and modern English Xu et al. (2012) (42K sentences), Amazon (558K) and Yelp (448K) sentiment datasets [^1] https://github.com/lijuncen/Sentiment-and-Style-Transfer (two styles each). We follow prior work by using the existing train-dev-test splits. In addition to these datasets, we use the StylePTB Lyu et al. (2021) dataset containing 21 individual and 32 compositional fine-grained style changes.
@@ 5 Conclusion
Learning generic, high-level text style representations from low-level, linguistically motivated changes and generalizing to high-level styles according to Assumption 1 by Lyu et al. (2021) presents challenges. Although the approach does not yield the expected results using contrastive learning, cross-entropy loss shows improvements for some settings. However, our approach does not yield the expected results compared to previous work (Patel et al., 2024) using contrastive learning to learn generic style representations.

### v2

@@ 1 Introduction
Learning meaningful representations of text has received much attention recently as models pre-trained on large corpora have become the standard for extracting representations that capture prior knowledge (i.a. Devlin et al., 2019; Radford and Narasimhan, 2018). However, most pre-trained models focus on general aspects, as they are trained with causal or masked language modeling objectives. Therefore, they are suboptimal for NLP tasks focusing on a specific aspect, such as style. There are many notable NLP tasks, such as identifying chatbot-written text based on its style (e.g., Soto et al. (2024)) or style transfer where models rely on dedicated style representations (e.g., John et al. (2019)).
@@
Based on the hypothesis that low-level style changes compose high-level style changes, this work explores learning generic, sentence-level style representations. We take a pre-trained encoder model producing general-purpose representations and fine-tune it to distinguish between low-level stylistic changes using contrastive learning and cross-entropy loss. We hypothesize that the resulting text encoder generalizes to high-level styles, applying the view by Lyu et al. (2021) on how low-level stylistic changes compose high-level changes to representation learning for styles.
@@ 3 Method
Assumption 2
Learning style embeddings can be achieved by contrasting different styles (Patel et al., 2024).
We combine the aforementioned assumptions to come up with the following method.
@@ 4.1 Experimental Setup
Fine-Tuning Data
We use the StylePTB Lyu et al. (2021) dataset containing 21 individual and 32 compositional fine-grained style changes to train our generic text style embeddings.
Evaluation Data
We evaluate the learned style embeddings on datasets that take a data-driven approach, containing high-level stylistic changes, to define a text style. These are used in many recent works on text style transfer. We experiment with the Bible corpus Carlson et al. (2018) with eight different styles of the Bible (249K sentences), Grammarly’s Yahoo Answers Formality Corpus (GYAFC) Rao and Tetreault (2018) (113K sentences) in two styles, a collection of Shakespeare plays in Shakespearean and modern English Xu et al. (2012) (42K sentences), Amazon (558K) and Yelp (448K) sentiment datasets [^1] https://github.com/lijuncen/Sentiment-and-Style-Transfer (two styles each). We follow prior work by using the existing train-dev-test splits.
@@ 5 Conclusion
Learning generic, high-level text style representations from low-level, linguistically motivated changes and generalizing to high-level styles according to Assumption 1 by Lyu et al. (2021) using contrastive learning according to Assumption 2 by Patel et al. (2024) presents challenges. Although the approach does not yield the expected results using contrastive learning, cross-entropy loss shows improvements for some settings. However, our approach does not yield the expected results compared to previous work (Patel et al., 2024) using contrastive learning to learn generic style representations.
@@
Acknowledgements
Part of this work was conducted within the DFG Research Unit FOR 5359 on Deep Learning on Sparse Chemical Process Data (BU 4042/2-1, KL 2698/6-1, and KL 2698/7-1) and the DFG TRR 375 (no. 511263698). SF and MK acknowledge support by the Carl-Zeiss Foundation (AI-Care, Process Engineering 4.0), the BMBF award 01|S2407A, and the DFG awards FE 2282/6-1, FE 2282/1-2, KL 2698/5-2, and KL 2698/11-1.